# Formatting Instructions For NeurIPS 2025

### v1

@@ Abstract
Internal representations are crucial for understanding deep neural networks, such as their properties and reasoning patterns, but remain difficult to interpret. While mapping from feature space to input space aids in interpreting the former, existing approaches often rely on crude approximations. We propose using a conditional diffusion model – a pretrained high-fidelity diffusion model conditioned on spatially resolved feature maps – to learn such a mapping in a probabilistic manner. We demonstrate the feasibility of this approach across various pretrained image classifiers from CNNs to ViTs, showing excellent reconstruction capabilities. Through qualitative comparisons and robustness analysis, we validate our method and showcase possible applications, such as the visualization of concept steering in input space or investigations of the composite nature of the feature space. This approach has broad potential for improving feature space understanding in computer vision models.
@@ 1 Introduction
The feature space is vital for understanding neural network decision processes as it offers insights into the internal representations formed by these models as they process input data. While it serves as the foundation for many modern explainability approaches [21, 3], its importance extends beyond interpretability. The feature space provides a rich resource for investigating fundamental properties of deep neural networks, including their robustness against perturbations, invariance characteristics, and symmetry properties [4]. By analyzing the geometry and topology of these learned representations, researchers can gain insights into model generalization capabilities, failure modes, and the emergence of higher-order patterns in the data. This perspective enables advancements in theoretical understanding of neural networks while informing practical improvements in architecture design and training methodologies.
An important challenge in examining the feature space is establishing a connection back to the input domain, especially for classification models that map to labels rather than the same domain as the input. One aspect of this challenge involves identifying which part of the input a particular region or unit in feature space is sensitive to. GradCAM [24] pioneered this by linearly upsampling a region of interest in feature space to the input size. However, linear upsampling imposes a rather strong implicit assumption. As an alternative, one might consider the entire receptive field of a feature map location, yet in deep architectures these fields tend to be broad and less informative.
The more intricate second aspect of this challenge is to derive a mapping from the entirety of the feature space representation back to the input domain – beyond mere localization. Recent works proposed to leverage conditional generative models to learn such a mapping by conditioning them on feature maps [4, 6, 23]. However, these approaches either build on pooled feature maps (discarding finegrained spatial details of the feature map), only provide deterministic mappings (overlooking the inherent uncertainty of this ill-posed problem), or do not utilize state-of-the-art generative models. To the best of our knowledge, there is no probabilistic model that provides high-fidelity input samples when conditioned on a spatially resolved feature map – thereby integrating both aspects of the challenge described above. We aim to close this gap with this submission.
@@
    We demonstrate the feasibility of learning high-fidelity mappings from feature space to input space using a conditional diffusion model of the ControlNet-flavor, as exemplified in Fig. 1. We investigate this for different computer vision models, ranging from CNNs to ViTs.
@@
    We provide quantitative evidence that generated samples align with the feature maps of the original samples and that the samples represent high-fidelity natural images, see Tab. 1. and carry out a qualitative model comparison, see Fig. 3 as well as a robustness analysis, see Tab. 2.
@@ 2 Methods
Approach In this work, we propose a method called FeatInv to approximate an inverse mapping from a model’s feature space to input space. Our method conditions a pretrained stable diffusion model on a spatially resolved feature map extracted from a pretrained CNN/ViT model of our choice. As described in detail in the next paragraph, the feature maps are provided as conditional information along with an unspecific text prompt (“a high-quality, detailed, and professional image”) to a conditional diffusion model of the ControlNet [29] flavor. Importantly, rather than achieving a precise reconstruction of the original sample in input space, our goal is to infer high-fidelity, synthetic images whose feature representations align with those of the original image when passed through a pretrained CNN/ViT model.
@@
Architecture and training procedure We use a ControlNet [29] architecture, building on a pretrained diffusion models, in our case a MiniSD [20] model operating at an input resolution of $256\times 256$. The ControlNet is a popular approach to condition a pretrained diffusion model on dense inputs such as segmentation maps or depth maps. It leverages a pretrained (text-conditional) diffusion model, whose weights are kept frozen. The trainable part of the ControlNet model mimics the internal structure of the pretrained diffusion model, with additional layers introduced to incorporate conditioning inputs. These conditional inputs are processed by a dedicated encoder and inserted into the corresponding computational blocks, where their outputs are added to those of the original diffusion model. Convolutional layer that are initialized to zero ensure that the optimization of the ControlNet model starts from the pretrained diffusion model.
@@
Conditional input encoder An import design choice is the conditional input encoder, which maps the feature map (with shape $H_{f}\times W_{f}\times C_{f}$, where $H_{f}$,$W_{f}$,$C_{f}$ correspond to the height, width and channels of the feature map, respectively) to the diffusion model’s internal representation space (with shape $H_{d}\times W_{d}\times C_{d}$). As a definite example for $224\times 224$ input resolution, for the output of ResNet50’s final convolutional block with $H_{f}=W_{f}=7$, $C_{f}=2048$ we learn a mapping to the diffusion model’s internal representation space with $H_{d}=W_{d}=32$ and $C_{d}=320$. To this end, we first use bilinear upsampling to reach the target resolution. Then, we allow for a shallow CNN to learn a suitable mapping from the model’s representation space to the diffusion model’s representation space.
@@
Pooled vs. unpooled To demonstrate superiority over prior work [4], we also consider the case of pooled feature representations obtained from average-pooling spatial tokens/feature maps. In order to process them using the same pipeline as for conditioning on spatially resolved feature maps, we copy the $C_{d}$-dimensional input vector along $H_{d}$ and $W_{d}$ times to reach an input tensor with shape $H_{d}\times W_{d}\times C_{d}$ as before.
@@
Training The ControlNet is trained using the same noise prediction objective as the original diffusion model [12]. Control signals are injected at multiple layers throughout the network, rather than being restricted to the middle layers, allowing them to influence the denoising process at various stages. Training was conducted on the ImageNet training set with a batch size of 8 and a learning rate of 1e-5 using an AdamW optimizer with the stable diffusion model locked. The ControlNet was trained on ImageNet for three epochs over approximately 45 to 60 hours (depending on backbone) of compute time on two NVIDIA L40 GPUs. During the course of this project, about five times more models were trained until the described setup was reached.
@@
Full pipeline We work with the original input resolution of the respective pretrained models, which varies between $224\times 224$ and $384\times 384$ for the considered models, see the supplementary material for a detailed breakdown. Even though the approach allows conditioning on any feature map, we restrict ourselves to the last spatially resolved feature map, i.e., directly before the pooling layer, and learn mappings to MiniSD’s internal feature space. The MiniSD model always returns an image with resolution $256\times 256$, which we upsample/downsample to the model’s expect input resolution via bilinear upsampling/downsampling. The full generation pipeline is visualized in Fig. 2.
@@ 3 Related Work
Conditional diffusion models Achieving spatially controllable image generation while leveraging a pretrained diffusion model has been a very active area of research recently, see [30] for a recent review. Applications include the conditional generation of images from depth maps, normal maps or canny maps. Popular approaches in this direction include ControlNet [29] or GLIGEN [14]. The mapping from feature maps as conditional input is structurally similar to the mentioned cases of spatially controllable generation. However, there is a key distinction. In the previously mentioned cases, the conditional information typically matches the resolution of the input image. This often necessitates downsampling to reach the diffusion model’s internal representation space. In contrast, commonly used classification models (including CNNs and vision transformers) leverage feature maps with a reduced spatial resolution. Consequently, the spatial resolution of the conditional information is typically lower dimensional than the diffusion model’s internal representation space. This difference necessitates an upsampling operation before conditioning on feature maps.
@@
Feature visualization The idea to reveal structures in feature space to understand what a neural network has learned is an old one. Approaches range from identifying input structures or samples that maximize the activation of certain feature neurons [7, 19] to approximate inversion of the mapping from input to features space [28]. Our approach clearly stands in the tradition of the latter approach. Previous work has attempted to learn a deterministic mapping that “inverts” AlexNet feature maps [6]. This approach was recently extended to invert vision transformer representations [22]. In contrast, FeatInv learns a probabilistic mapping using state-of-the-art diffusion models and investigates state-of-the-art model architectures. Other approaches tackle the problem using invertible neural networks to connect VAE latent representations to input space [23] and/or disentangle these representations using concept supervision [8]. In contrast, FeatInv does not rely on a particular encoder/decoder structure but can use any pretrained neural network as encoder. The closest prior work to our approach is [4], which also uses a diffusion model to learn a mapping from feature space to input space. However, it uses pooled representations as input, i.e. neglects the spatial resolution of the feature map. We argue that pooled representations are too coarse for many applications as they disregard the finegrained spatial structure of the feature space.
@@
Representation surgery Finally, related feature inversion approaches have also been explored beyond computer vision, for example in natural language processing [18]. Here, the ability to invert latent representations is seen as an essential component for representation surgery approaches [1]. FeatInv enables similar approaches for computer vision models.
@@
Table 1: Reconstruction quality and image quality of the individual models: For the three considered backbones, we indicate three performance metrics to assess the reconstruction quality: Cosine similarity in feature space (cosine-sim), calculated by taking the median of the cosine similarity of all superpixels, top5(1) matches using the top1 prediction of the original sample as ground truth (top5(1) match) and FID-scores (FID) to assess the quality of the generated samples. We consider generative models conditioned on unpooled feature maps (rows 1-3) and models conditioned on pooled feature maps (rows 4-6). The results indicate that the proposed approach produces high-fidelity input samples as perceived by the respective models.
@@
unpooled ResNet50 0.40 87% (63%) 12.32
ConvNeXt 0.61 94% (76%) 7.14
SwinV2 0.57 95% (81%) 11.56
pooled ResNet50 0.09 47% (22%) 26.37
ConvNeXt 0.18 42% (18%) 31.01
SwinV2 0.14 47% (23%) 36.21
4 Results
We investigate three models ResNet50 [11] (original torchvision weights), ConvNeXt [16] and SwinV2 [15] 111timm model weights: convnext_base.fb_in22k_ft_in1k, swinv2_base_window12to24_192to384_22kft1k. all of which have been pretrained/finetuned on ImageNet1k. ConvNeXt and SwinV2 represent modern convolution-based and vision-transformer-based architectures, identified as strong backbones in [10]. We include ResNet50 due to its widespread adoption. For each model, we train a conditional diffusion model conditioned on the representations of the last hidden layer before the final pooling layer to reconstruct the original input samples. Below, we report on quantitative and qualitative aspects of our findings.
4.1 Quantitative and qualitative comparison
Experimental setup For each ImageNet class, we reconstructed 10 validation set samples with FeatInv, resulting in 10000 reconstructed samples. We adjust the diffusion model’s guidance scale to optimize the FID score between original and reconstructed samples on the validation set, resulting in a guidance scale of 8(1.75) for (un)pooled conditional input, see the supplementary material for details. We reconstruct with a sample step size of 50. For each model this took roughly 12 hours on a single NVIDIA L40 GPU. The generated samples are assessed according to two complementary quality criteria, reconstruction quality and sample quality:
  1. 1.
    Reconstruction quality The encoded generated image should end up close to the feature representation of the original samples, which can be understood as a reconstruction objective that is implemented implicitly by conditioning the diffusion model on a chosen feature map. (1a) The most obvious metric is cosine similarity between both feature maps. However, not all parts of the feature space will be equally important for the downstream classifier. (1b) Most reliable measure is the classifier output itself. Focusing on top-predictions, one can also compare top-k predictions to the top prediction for the original sample. More general alignment measures between generated input and original feature representation are not helpful in this context, as we require a precise reconstruction of the original feature space for the downstream classifier above the layer under consideration.
  2. 2.
    Sample quality We aim to generate samples within the set of high-fidelity natural images. In our case, this objective is is implemented through the use of a pretrained diffusion model. Apart from qualitative assessments in the following sections, we rely on FID-scores as established measures to assess sample quality.
@@ 4.1 Quantitative and qualitative comparison
Reconstruction quality Comparing identical models conditioned either on pooled or unpooled feature maps, not surprisingly unpooled models show a significantly higher reconstruction quality. Samples generated by models conditioned on unpooled feature maps show a very good alignment with the feature maps of the original samples (cosine similarities above 0.57 and top5 matching predictions of 94% or higher for the two modern vision backbones). Samples conditioned on pooled feature maps show some alignment but fail to accurately reconstruct the respective feature map and are therefore unreliable for investigations of structural properties of models. These findings support the hypothesis that the approach yield feature space reconstructions that closely match the original feature representations.
@@
Sample quality The corresponding class-dependent diffusion model achieves an FID score around 29, which is typically considered as good quality. The models conditioned on pooled representations still show acceptable FID scores between 26 and 36. Interestingly, models conditioned on unpooled representations show a significant increase in image quality with FID scores between 7 and 12. These results support the statement that the created samples were sampled from the space of high-fidelity natural images.
@@
Backbone comparison Within each category (pooled vs. unpooled), there is a gap between the two most recent model architectures ConvNeXt and SwinV2, notwithstanding the architectural differences (CNN vs. Vision transformer) between the two, in comparison to the older Resnet50 models. The former achieve cosine similarities of .6 or higher and top5 matches of 95% or higher in the unpooled category. This suggests that there is a qualitative difference between the representations of ResNet50-representations and representations of more modern image backbones.
Figure 3: Qualitative comparison of reconstructed samples for the ResNet50, ConvNeXt, SwinV2 and SwinV2 pooled models. The cosine similarity of the original feature map and that of the reconstruction is noted at the bottom edge of the images. The qualitative comparison confirms the insights from the quantitative analysis in Tab. 1. The two modern vision backbones, ConvNeXt and SwinV2, show reconstructions that resemble the original very closely, not only in terms of semantic content and spatial alignment but also in terms of color schemes and finegrained details. Semantic content and composition also mostly matches in case of the ResNet50, but not even the semantic content seems to be captured when using pooled representations (SwinV2 pooled as an example).
@@
Qualitative comparison In Fig. 3, we present a qualitative comparison based on randomly selected samples. The visual impressions of ConvNeXt and SwinV2 reconstructions are similar to each other while also being close to the input sample despite the fact that they were trained on high-level semantic feature maps, i.e., without a reconstruction objective in input space. The ResNet50 reconstructions seem in many cases an interpretation of the sample’s semantic content (see e.g. 2. toucan or 5. file), albeit with the correct spatial composition, while matching specific color composition and textures much less accurately than ConvNeXt and SwinV2. We primarily attribute the differences between ResNet and ConvNeXt/SwinV2 to the nature of the feature spaces themselves, stressing qualitative difference between modern architectures such as ConvNeXt and SwinV2 and older model architectures such as ResNet50, which are much more pronounced than the differences between different model architectures such as ViTs and CNNs. The samples obtained from conditioning on pooled feature representations often seem to capture overall semantic content of the image correctly (file, space shuttle, traffic light), but fail to reflect the details of the composition of the image.

### v2

# FeatInv: Spatially resolved mapping from feature space to input space using conditional diffusion models

@@ Abstract
Internal representations are crucial for understanding deep neural networks, such as their properties and reasoning patterns, but remain difficult to interpret. While mapping from feature space to input space aids in interpreting the former, existing approaches often rely on crude approximations. We propose using a conditional diffusion model - a pretrained high-fidelity diffusion model conditioned on spatially resolved feature maps - to learn such a mapping in a probabilistic manner. We demonstrate the feasibility of this approach across various pretrained image classifiers from CNNs to ViTs, showing excellent reconstruction capabilities. Through qualitative comparisons and robustness analysis, we validate our method and showcase possible applications, such as the visualization of concept steering in input space or investigations of the composite nature of the feature space. This approach has broad potential for improving feature space understanding in computer vision models.
@@ 1 Introduction
The feature space is vital for understanding neural network decision processes as it offers insights into the internal representations formed by these models as they process input data. While it serves as the foundation for many modern explainability approaches (Rai et al., 2024; Bereska & Gavves, 2024), its importance extends beyond interpretability. The feature space provides a rich resource for investigating fundamental properties of deep neural networks, including their robustness against perturbations, invariance characteristics, and symmetry properties (Bordes et al., 2022). By analyzing the geometry and topology of these learned representations, researchers can gain insights into model generalization capabilities, failure modes, and the emergence of higher-order patterns in the data. This perspective enables advancements in theoretical understanding of neural networks while informing practical improvements in architecture design and training methodologies.
An important challenge in examining the feature space is establishing a connection back to the input domain, especially for classification models that map to labels rather than the same domain as the input. One aspect of this challenge involves identifying which part of the input a particular region or unit in feature space is sensitive to. GradCAM (Selvaraju et al., 2017) pioneered this by linearly upsampling a region of interest in feature space to the input size. However, linear upsampling imposes a rather strong implicit assumption. As an alternative, one might consider the entire receptive field of a feature map location, yet in deep architectures these fields tend to be broad and less informative.
The more intricate second aspect of this challenge is to derive a mapping from the entirety of the feature space representation back to the input domain – beyond mere localization. Recent works proposed to leverage conditional generative models to learn such a mapping by conditioning them on feature maps (Bordes et al., 2022; Dosovitskiy & Brox, 2016; Rombach et al., 2020). However, these approaches either build on pooled feature maps (discarding finegrained spatial details of the feature map), only provide deterministic mappings (overlooking the inherent uncertainty of this ill-posed problem), or do not utilize state-of-the-art generative models. Related approaches such as diffusion autoencoders (Preechakul et al., 2022) show that diffusion models can indeed be fitted with meaningful and decodable latent representations that enable near-exact reconstructions and the manipulation of semantic attributes. However, their latent codes are global and not spatial, whereas our focus is on conditioning spatially resolved feature maps to preserve the fine-grained structure. We argue that a lack of faithfulness, i.e., the inability to reconstruct spatially resolved feature maps, is detrimental for downstream applications, e.g., in explainable AI or for robustness analyses, as it fails to reflect the true relation between feature space and input space in the model. In addition, pooled representations are not meaningful for the purpose of studying feature representations of intermediate layers. To the best of our knowledge, there is no probabilistic model that provides high-fidelity input samples when conditioned on a spatially resolved feature map – thereby integrating both aspects of the challenge described above. We aim to close this gap with this submission.
@@
    We demonstrate the feasibility of learning high-fidelity mappings from a spatially resolved feature space to input space using a conditional diffusion model of the ControlNet-flavor, as exemplified in Fig. 1. We investigate this for different computer vision models, ranging from CNNs to ViTs.
@@
    We provide quantitative evidence that generated samples align with the feature maps of the original samples and that the samples represent high-fidelity natural images, see Tab. 1. and carry out a qualitative model comparison, see Fig. 3.
@@
    We provide robustness analysis, see Tab. 2 as well as further analysis demonstrating the generalizability of our approach across different ConvNeXt stages (see Fig. 4), its behavior on out-of-distribution data (see Fig. 5), as well as its robustness to corrupted images (see Sec. 4.2).
  4. 4.
@@ 2 Methods
Approach In this work, we propose a method called FeatInv to approximate an inverse mapping from a model’s feature space to input space. Our method conditions a pretrained stable diffusion model on a spatially resolved feature map extracted from a pretrained CNN/ViT model of our choice. As described in detail in the next paragraph, the feature maps are provided as conditional information along with an unspecific text prompt (“a high-quality, detailed, and professional image”) to a conditional diffusion model of the ControlNet (Zhang et al., 2023a) flavor. Importantly, rather than achieving a precise reconstruction of the original sample in input space, our goal is to infer high-fidelity, synthetic images whose feature representations align with those of the original image when passed through a pretrained CNN/ViT model.
@@
Architecture and training procedure We use a ControlNet (Zhang et al., 2023a) architecture, building on a pretrained diffusion models, in our case a MiniSD (Pinkney, 2023) model operating at an input resolution of $256\times 256$. The ControlNet is a popular approach to condition a pretrained diffusion model on dense inputs such as segmentation maps or depth maps. It leverages a pretrained (text-conditional) diffusion model, whose weights are kept frozen. The trainable part of the ControlNet model mimics the internal structure of the pretrained diffusion model, with additional layers introduced to incorporate conditioning inputs. These conditional inputs are processed by a dedicated encoder and inserted into the corresponding computational blocks, where their outputs are added to those of the original diffusion model. Convolutional layers initialised to zero ensure that optimisation of the ControlNet model starts from the pretrained diffusion model.
@@
Conditional input encoder An important design choice is the conditional input encoder, which maps the feature map (with shape $H_{f}\times W_{f}\times C_{f}$, where $H_{f}$$W_{f}$$C_{f}$ correspond to the height, width and channels of the feature map, respectively) to the diffusion model’s internal representation space (with shape $H_{d}\times W_{d}\times C_{d}$). As a definite example for $224\times 224$ input resolution, for the output of ResNet50’s final convolutional block with $H_{f}=W_{f}=7$, $C_{f}=2048$ we learn a mapping to the diffusion model’s internal representation space with $H_{d}=W_{d}=32$ and $C_{d}=320$. To this end, we first use bilinear upsampling to reach the target resolution. Then, we allow for a shallow CNN to learn a suitable mapping from the model’s representation space to the diffusion model’s representation space.
@@
Pooled vs. unpooled To demonstrate superiority over prior work (Bordes et al., 2022), we also consider the case of pooled feature representations obtained from average-pooling spatial tokens/feature maps. In order to process them using the same pipeline as for conditioning on spatially resolved feature maps, we copy the $C_{d}$-dimensional input vector along $H_{d}$ and $W_{d}$ times to reach an input tensor with shape $H_{d}\times W_{d}\times C_{d}$ as before.
@@
Training The ControlNet is trained using the same noise prediction objective as the original diffusion model (Ho et al., 2020). Control signals are injected at multiple layers throughout the network, rather than being restricted to the middle layers, allowing them to influence the denoising process at various stages. Training was conducted on the ImageNet training set with a batch size of 8 and a learning rate of 1e-5 using an AdamW optimizer with the stable diffusion model locked. The ControlNet was trained on ImageNet for three epochs over approximately 45 to 60 hours (depending on backbone) of compute time on two NVIDIA L40 GPUs. During the course of this project, about five times more models were trained until the described setup was reached.
@@
Full pipeline We work with the original input resolution of the respective pretrained models, which varies between $224\times 224$ and $384\times 384$ for the considered models, see the Supplementary Material A.1 for a detailed breakdown. Even though the approach allows conditioning on any feature map, we restrict ourselves to the last spatially resolved feature map, i.e., directly before the pooling layer, and learn mappings to MiniSD’s internal feature space. The MiniSD model always returns an image with resolution $256\times 256$, which we upsample/downsample to the model’s expect input resolution via bilinear upsampling/downsampling. The full generation pipeline is visualized in Fig. 2.
@@ 3 Related Work
Conditional diffusion models Achieving spatially controllable image generation while leveraging a pretrained diffusion model has been a very active area of research recently, see (Zhang et al., 2023b) for a recent review. Applications include the conditional generation of images from depth maps, normal maps or canny maps. Popular approaches in this direction include ControlNet (Zhang et al., 2023a) or GLIGEN (Li et al., 2023). The mapping from feature maps as conditional input is structurally similar to the mentioned cases of spatially controllable generation. However, there is a key distinction. In the previously mentioned cases, the conditional information typically matches the resolution of the input image. This often necessitates downsampling to reach the diffusion model’s internal representation space. In contrast, commonly used classification models (including CNNs and vision transformers) leverage feature maps with a reduced spatial resolution. Consequently, the spatial resolution of the conditional information is typically lower dimensional than the diffusion model’s internal representation space. This difference necessitates an upsampling operation before conditioning on feature maps.
@@
Feature visualization The idea to reveal structures in feature space to understand what a neural network has learned is an old one. Approaches range from identifying input structures or samples that maximize the activation of certain feature neurons (Erhan et al., 2009; Nguyen et al., 2016) to approximate inversion of the mapping from input to features space (Zeiler, 2014). Our approach clearly stands in the tradition of the latter approach. Previous work has attempted to learn a deterministic mapping that “inverts” AlexNet feature maps (Dosovitskiy & Brox, 2016). This approach was recently extended to invert vision transformer representations (Rathjens et al., 2024). In contrast, FeatInv learns a probabilistic mapping using state-of-the-art diffusion models and investigates state-of-the-art model architectures. Other approaches tackle the problem using invertible neural networks to connect VAE latent representations to input space (Rombach et al., 2020) and/or disentangle these representations using concept supervision (Esser et al., 2020). In contrast, FeatInv does not rely on a particular encoder/decoder structure but can use any pretrained neural network as encoder. The closest prior work to our approach is (Bordes et al., 2022), which also uses a diffusion model to learn a mapping from feature space to input space. However, it uses pooled representations as input, i.e. neglects the spatial resolution of the feature map. We argue that pooled representations are too coarse for many applications as they disregard the finegrained spatial structure of the feature space. Diffusion autoencoders (Preechakul et al., 2022) also explore how diffusion-based representations can be made both meaningful and decodable, but their latent codes are global vectors. In contrast, we condition directly on spatial feature maps to preserve fine-grained structure.
@@
Representation surgery Finally, related feature inversion approaches have also been explored beyond computer vision, for example in natural language processing (Morris et al., 2023). Here, the ability to invert latent representations is seen as an essential component for representation surgery approaches (Avitan et al., 2025). FeatInv enables similar approaches for computer vision models.
4 Results
We investigate three models ResNet50 (He et al., 2016) (original torchvision weights), ConvNeXt (Liu et al., 2022b) and SwinV2 (Liu et al., 2022a) 111timm model weights: convnext_base.fb_in22k_ft_in1k, swinv2_base_window12to24_192to384_22kft1k. all of which have been pretrained/finetuned on ImageNet1k. ConvNeXt and SwinV2 represent modern convolution-based and vision-transformer-based architectures, identified as strong backbones in (Goldblum et al., 2024). We include ResNet50 due to its widespread adoption. For each model, we train a conditional diffusion model conditioned on the representations of the last hidden layer before the final pooling layer to reconstruct the original input samples. Below, we report on quantitative and qualitative aspects of our findings.
4.1 Quantitative and qualitative comparison
Experimental setup For each ImageNet class, we reconstructed 10 validation set samples with FeatInv, resulting in 10,000 reconstructed samples. We adjust the diffusion model’s control strength and guidance scale to optimize the match of classification outputs between original and reconstructed samples on the validation set, resulting in different control strengths and guidance scales for each model. Since our goal is to represent the feature space as accurately as possible, we also observed the cosine similarity between the feature maps of the original and reconstructed samples and observed a strong correlation between classification match and cosine similarity, which further supports our choice of parameters. In our experience, it is also possible to achieve good results with the same control strength and guidance scale for all three models, see the Supplementary Material A.2 for details. We reconstruct with a sample step size of 50. For each model this took roughly 12 hours on a single NVIDIA L40 GPU. We only generate one sample per feature map but it is also possible to generate multiple to observe the variability across reconstructions (given the same conditional input), see Supplementary Material A.3 for insights. The generated samples are assessed according to two complementary quality criteria, reconstruction quality and sample quality:
  1. 1.
    Reconstruction quality The encoded generated image should end up close to the feature representation of the original samples, which can be understood as a reconstruction objective that is implemented implicitly by conditioning the diffusion model on a chosen feature map. (1a) The most obvious metric is cosine similarity between both feature maps. However, not all parts of the feature space will be equally important for the downstream classifier. (1b) Most reliable measure is the classifier output itself. Focusing on top-predictions, one can also compare top-k predictions to the top prediction for the original sample. More general alignment measures between generated input and original feature representation are not helpful in this context, as we require a precise reconstruction of the original feature space for the downstream classifier above the layer under consideration. A feature inversion method is deemed faithful if the reconstructed samples exhibit strong correspondence with the original samples, as assessed by the two metrics introduced above.
  2. 2.
    Sample quality We aim to generate samples within the set of high-fidelity natural images. In our case, this objective is implemented through the use of a pretrained diffusion model. Apart from qualitative assessments in the following sections, we rely on FID-scores as established measures to assess sample quality.
@@ 4.1 Quantitative and qualitative comparison
Table 1: Reconstruction quality and image quality of the individual models: For the three considered backbones, we indicate three performance metrics to assess the quality of our reconstructions: Cosine similarity in feature space (cosine-sim), calculated by averaging the cosine similarity of all superpixels while maintaining the dimensions of the original input (288 x 288) and upscaling the reconstruction (from 256 x 256), top5(1) matches using the top-1 prediction of the original sample as ground truth (top5(1) match), and FID scores (FID) to assess the quality of the generated samples. We consider generative models conditioned on unpooled feature maps (rows 1-3) and models conditioned on pooled feature maps (rows 4-6). The results indicate that the proposed approach produces faithful reconstructions with high fidelity that are both perceptually realistic and, when re-encoded, have similar values in the feature space as the original input.
@@
unpooled ResNet50 0.46 91% (70%) 11.49
ConvNeXt 0.61 94% (77%) 8.20
SwinV2 0.53 95% (80%) 12.69
pooled ResNet50 0.12 48% (23%) 31.64
ConvNeXt 0.19 44% (20%) 31.67
SwinV2 0.16 47% (22%) 49.04
@@
Reconstruction quality Comparing identical models conditioned either on pooled or unpooled feature maps, not surprisingly unpooled models show a significantly higher reconstruction quality. Samples generated by models conditioned on unpooled feature maps show a very good alignment with the feature maps of the original samples (cosine similarities above 0.53 and top5 matching predictions of 94% or higher for the two modern vision backbones). Samples conditioned on pooled feature maps show some alignment but fail to accurately reconstruct the respective feature map and are therefore unreliable for investigations of structural properties of models. These findings support the hypothesis that the approach yield feature space reconstructions that closely match the original feature representations. We also directly compared our pooled results to the approach proposed by Bordes et al. (2022), see Supplementary Material A.5 for details.
@@
Sample quality The corresponding class-dependent diffusion model achieves an FID score around 29, which is typically considered as good quality. The models conditioned on pooled representations still show acceptable FID scores between 31 and 49. Interestingly, models conditioned on unpooled representations show a significant increase in image quality with FID scores between 8 and 12. These results support the statement that the created samples were sampled from the space of high-fidelity natural images.
@@
Backbone comparison Within each category (pooled vs. unpooled), there is a gap between the two most recent model architectures ConvNeXt and SwinV2, notwithstanding the architectural differences (CNN vs. Vision transformer) between the two, in comparison to the older Resnet50 models. The former achieve cosine similarities of .61 or higher and top5 matches of 94% or higher in the unpooled category. This suggests that there is a qualitative difference between the representations of ResNet50-representations and representations of more modern image backbones.
@@
Qualitative comparison In Fig. 3, we present a qualitative comparison based on randomly selected samples. The visual impressions of ConvNeXt and SwinV2 reconstructions are similar to each other while also being close to the input sample despite the fact that they were trained on high-level semantic feature maps, i.e., without a reconstruction objective in input space. The ResNet50 reconstructions seem in many cases an interpretation of the sample’s semantic content (see e.g. 2. toucan or 5. file), albeit with the correct spatial composition, while matching specific color composition and textures much less accurately than ConvNeXt and SwinV2. We primarily attribute the differences between ResNet and ConvNeXt/SwinV2 to the nature of the feature spaces themselves, stressing qualitative difference between modern architectures such as ConvNeXt and SwinV2 and older model architectures such as ResNet50, which are much more pronounced than the differences between different model architectures such as ViTs and CNNs. The samples obtained from conditioning on pooled feature representations often seem to capture overall semantic content of the image correctly (file, space shuttle, traffic light), but fail to reflect the details of the composition of the image. This can further be observed in the Supplementary Material A.4.
Figure 3: Qualitative comparison of reconstructed samples for the ResNet50, ConvNeXt, SwinV2 and ConvNeXt pooled models. The cosine similarity of the original feature map and that of the reconstruction is noted at the bottom edge of the images. The qualitative comparison confirms the insights from the quantitative analysis in Tab. 1. The two modern vision backbones, ConvNeXt and SwinV2, show reconstructions that resemble the original very closely, not only in terms of semantic content and spatial alignment but also in terms of color schemes and finegrained details. Semantic content and composition also mostly matches in case of the ResNet50, but not even the semantic content seems to be captured when using pooled representations (ConvNeXt pooled as an example).
Conditioning on intermediate feature layers The experiments so far focused on the last feature layer before the classifier in order to be able to visualize the features that ultimately lead to the classification decision (e.g., for FeatInv-Viz in Section 4.3). However, FeatInv  can also applied to intermediate feature layers. For demonstration, we trained three additional FeatInv  models conditioned on the respective last feature of the first, second, and third stage within the ConvNeXt model. As expected, conditioning on earlier feature provides more specific spatial context and therefore leads to more accurate and detailed reconstructions, see the qualitative examples in Figure 14 in the supplemental material. As in Table 1, we can also assess the reconstruction quality and the image quality quantitatively. The main difference compared to Table 1 is the ability to consider not only the cosine similarity compared to the feature layer but also compared to other feature layers, which were here taken identical to the three intermediate layers used for conditioning above. The results of this experiment are compiled in Figure 4. Most notably, the faithfulness extends beyond the layer that was used for conditioning, i.e., the models generally achieve good reconstruction across all considered feature layers. In line with expectations, the similarity increases when conditioning on earlier feature layers as they provide more finegrained information on the composition of the image. Furthermore, it is interesting to note that the similarity increases consistently for all models with the later stages the feature maps were extracted from (except the last one). This suggests that the last feature maps from the second and third stages are less sensitive to reconstruction-related pixel deviations and instead represent more robust, semantically stable image features.
(a) cosine-similarity
(b) FID score
(c) top5(1) matching
Figure 4: Comparison of the cosine similarity, FID score and top5(1)match for the models trained on the last feature maps of the first, second, and third stage of ConvNeXt calculated by averaging the cosine similarity of all superpixels. The models trained on earlier features and reconstructing them perform better in terms of cosine-sim, FID score and matchings than the one which was trained on the very last feature map
4.2 Robustness evaluation
Cross-model evaluation To assess the robustness of the presented results, we carry out cross-model comparisons where we measure model performance based on samples generated by conditioning on the feature map extract from a different model. The results for this experiment are compiled in Tab. 2. It turns out that all three sets of samples (conditioned on features generated by the three different backbones) transfer quite remarkably to other models.